OpenStudy (anonymous):

A rabbit's fur grows at the rate of 5 over 2 inches per month. How many months will it take the rabbit's fur to grow 15 over 4 inches? please help

OpenStudy (anonymous):

you any answer

OpenStudy (anonymous):

hi welcome to openstudy @Hanaabdalah :) let y be the length of the fur each month and let x be the number of months. so it will be: y = 5/2 x Solve for x when y= 15/4. 15/4 = 5/2 x , can you solve x from here?

OpenStudy (anonymous):

i didn't know it

OpenStudy (is3535):

Month and a half.

OpenStudy (is3535):

medal

OpenStudy (anonymous):

to solve for x, divide both sides by 5/2 can you do this step? @is3535 can you explain how you did it? or just show it to the asker? :)

OpenStudy (anonymous):

i don't know it

OpenStudy (is3535):

i dont have no money

OpenStudy (anonymous):

what??

OpenStudy (is3535):

Month and a half. that your answer

OpenStudy (anonymous):

@Hanaabdalah I'm teaching you how to do it. What will you get when you divide 15/4 by 5/2? @is3535 please don't give direct answer because no one will learn from it and it is against our code of conduct. Thanks!

OpenStudy (is3535):

srry can i have a medal

OpenStudy (anonymous):

is it 1 month and 1/2